Cosec theta-cot theta=p,find the value of cosec theta?

Grade:10

2 Answers

Arun
6 years ago
Dear student
 
Cosec@ - cot@ = p
1 - cos@= p sin@
1- p sin@ = cos@
1- p sin@ = sqrt(1 - sin²@)
 
On square
1 + p² sin²@ - 2p sin@ = 1 - sin²@
Now 
You can make quadratic equation in sin@
Hence can find sin@ and cosec@
aftab
6 years ago
plz solve this question
a radioactive element decompose at the rate of 15% per hours.if a sample contain 10g of the element how much will be left after 2 hours?